# Cosmological perturbations of a relativistic MOND theory

### v1

@@ Abstract
A relativistic MOND theory, promising in reproducing cosmology as well as the MOND phenomenology in the low acceleration regime, was recently proposed. We present the post-Newtonian (PN) approximation and relativistic perturbation equations of this theory in cosmological context. The PN equations are presented to 1PN order and perturbation equations are presented in fully-nonlinear and exact forms. The gauge issues are clarified. The 1PN equations and linear perturbation equations are presented without imposing temporal gauge conditions. The MOND field can be interpreted as a fluid with specific equation of state without anisotropic stress, and the Jeans criterion is derived for the MOND field.
@@ I Introduction
Milgrom, in 1983, proposed MOND (Modified Newtonian Dynamics) to explain the gravity anomaly in galaxy rotations MOND-1983a . Instead of a popular way of saving conventional gravity (both Newton’s and Einstein’s) by introducing unseen dark matter (DM), he has modified in a non-relativistic context either gravity or inertia for the anomaly. Despite the lack of agreed relativistic theory, thus unable to address the cosmological demands other than the galactic rotation and related phenomena on similar scales, the non-relativistic proposal has survived till today as an alternative to the dark matter hypothesis. Recently, it gained independent observational support and challenge in stellar and Solar system scale phenomena; see Chae-2024 ; Hernandez-2024 for support from the long period binary stars observed by GAIA, and Vokrouhlicky-2024 ; Desmond-2024 for against in explaining dynamics and gravitational field in the solar system.
@@
Observations have pin down the Milgrom’s universal constant of nature $a_{M}\simeq 1.2\times 10^{-8}cm/sec^{2}$ McGaugh-2004 . As originally noticed by Milgrom MOND-1983a the transition acceleration $a_{M}$ is intriguingly close to the acceleration scale of the expanding Universe $cH_{0}/(2\pi)=1.1\times 10^{-8}cm/sec^{2}$ using the current value of Lemaître-Hubble parameter $H_{0}=70(km/sec)/Mpc$. The interpolating function $f$ is not determined yet.
Bekenstein and Milgrom, in 1984, proposed a non-relativistic Lagrangian formulation of MOND, often called as AQUAL, leading to a modified Poisson’s equation $\nabla\cdot(f\nabla\Phi)=4\pi G\varrho$ with $\vec{a}=\nabla\Phi$ Bekenstein-Milgrom-1984 . At large distance from an isolated body with high symmetry, Milgrom’s original proposal is recovered Bekenstein-Milgrom-1984 . In the MOND based on the modified Poisson’s equation, the internal gravitational dynamics of a system is affected by an external gravitational field, often called the external-field effect, thus violating the strong equivalence principle Bekenstein-Milgrom-1984 . Variety of modifications of theory producing the MOND phenomenology is suggested Milgrom-2010-QUMOND ; Milgrom-2023-TriMOND ; Milgrom-2023-GQUMOND , and diverse relativistic extensions are also made Bekenstein-Milgrom-1984 ; Famaey-2012 .
Blanchet and Marsat Blanchet-Marsat-2011 introduced a small addition of a field in Einstein’s gravity to recover the MOND phenomenology in weak gravity low acceleration regime. They used a function of the acceleration of a four-vector $U_{a}$ associated with a scalar function $\tau$ as $U_{a}\propto\nabla_{a}\tau$. The theory is motivated by a quantum gravity proposal with local Lorentz invariance violation Horava-2009 ; Blas-2010 and Einstein aether theories Jacobson-Mattingly-2001 ; Jacobson-2010 ; Sanders-2011 .
The MOND proposal is successful in explaining the phenomenology in the galactic scale without demanding DM. However, successful cosmology also demands the presence of a sort of DM (a fluid with dust-like equation of state) in explaining phenomena in the large-scale structure, the cosmic background radiation (CMB), and the background matter content. Blanchet and Skordis Blanchet-Skordis-2024 recently proposed one more additional function in Blanchet and Marsat’s model, now using a function of the normalization amplitude of the same four-vector $U_{a}$ to accommodate such demands in cosmology.
@@
Here, we present equations for the PN approximation and relativistic perturbation theory in cosmological context. The PN approximation is presented to 1PN order, and the perturbation theory is presented in fully-nonlinear and exact form. We address the gauge issues in both formulations and the 1PN equations and linear perturbation equations are presented without imposing temporal gauge conditions. These subjects are already studied in Blanchet-Skordis-2024 and here we elaborate some aspects. For both formulations, we compare the MOND contribution with an ordinary fluid. The MOND field does not contribute to the vector- and tensor-type perturbations, and it can be interpreted as a fluid with a specific equation of state (depending on the functional forms of ${\cal J}$ and ${\cal K}$) and is not associated with the anisotropic stress.
@@ II Relativistic MOND
$\hskip-22.76228ptU_{a}\equiv-{c\over Q}\nabla_{a}\tau,\quad Q\equiv c\sqrt{-(\nabla^{a}\tau)\nabla_{a}\tau},$
@@
and $L_{m}$ is the ordinary matter (baryon and photon) part Lagrangian. Although we call the fundamental field simply a $\tau$-field, what is introduced is a four-vector field $U_{a}$ defined using the $\tau$-field; $U_{a}$ is a normalized timelike four-vector with $U^{c}U_{c}\equiv-1$; $q_{ab}\equiv g_{ab}+U_{a}U_{b}$ is a spatial projection tensor orthogonal to $U_{a}$; $A_{a}$ is an acceleration of the four-vector. For $\tau=\tau(t)$, we have $U_{a}=n_{a}$ where $n_{a}$ is the normal four-vector with $n_{i}\equiv 0$.
@@
We will often consider two models for ${\cal K}$ Blanchet-Skordis-2024
@@ III Cosmological Background
where $x^{0}=\eta$ is the conformal time; $a$ is the cosmic scale factor. Using $\tau=\tau(t)$, we have
@@
where an overdot indicates a derivative with respect to the cosmic time $t$. Equations (4), with $T_{00}=a^{2}\mu$ and $T_{ij}=a^{2}p\delta_{ij}$, gives $\tau$-fluid quantities
@@ III.1 Background constraint
$\hskip-14.22636ptQ=1+\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n},$
$\hskip-14.22636pt\varrho_{\tau}={c^{2}I_{0}\over 8\pi Ga^{3}}\Big{[}1+{n\over n+1}\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]},$
$\hskip-14.22636ptw_{\tau}=\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{[}n+1+n\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]}^{-1},$
$\hskip-14.22636ptc_{\tau}^{2}={1\over n}\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{[}1+\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]}^{-1},$ (19)
where $a_{0}=a(t_{0})\equiv 1$ at present epoch; we used $\mu=\varrho c^{2}$. By a suitable parameter choice, with $[I_{0}/(2\nu^{2}{\cal K}_{n+1}a^{3})]^{1/n}\ll 1$, ${\cal K}$-term can simulate a dust fluid with $w_{\tau}=c_{\tau}^{2}=0$ in recent era. In the early era it approaches the stiff fluid with $w_{\tau}=c_{\tau}^{2}=1$ Blanchet-Skordis-2024 .
For the dust-like behavior of $\tau$-field in the background, the authors of Blanchet-Skordis-2024 demand $w_{\tau*}\leq 0.0164$ at $a_{*}\sim 10^{-4.5}$. Using $I_{0}\simeq 8\pi G\varrho_{\tau 0}/c^{2}=3\Omega_{\tau 0}H_{0}^{2}/c^{2}$, we have
@@
For $n=1$, with $\Omega_{\tau 0}=0.26$, the upper limit gives $L_{*}=0.22kpc$. As $n$ increases the upper limit increases: we have $L_{*}=0.34$, $0.46$, $0.58$, $0.71$, and $1.4kpc$, respectively, for $n=2,3,4,5$, and $10$. This is a constraint on $\nu$-parameter from the requirement of dust-like behavior of $\tau$-field in the background cosmology Blanchet-Skordis-2024 .
@@ IV PN approximation
where the index of $P_{i}$ is associated with $\delta_{ij}$ and its inverse. In this form we already imposed spatial gauge conditions Chandrasekhar-1965 ; Hwang-Noh-Puetzfeld-2008 . In PN approximation, we use $x^{0}=ct$. To 0PN order, $\Phi$ is the Newtonian gravitational potential; $\Psi$ and $P_{i}$ are 1PN order. To 1PN order, $\Phi$ in Eq. (59) contains 1PN order contribution as well. In Chandrasekhar’s notation, we have Chandrasekhar-1965 ; Hwang-Noh-Puetzfeld-2008
@@
where $\sigma$ and $\Sigma$ are functions of spacetime corresponding to 0PN and 1PN order, respectively. As we take $\tau=t$ to the background order, we have $Q=1$ thus ${\cal K}=0$, and the $\tau$-field does not contribute to the background order equations.
@@
Apparently, the expansion is lengthy if we keep perturbations of $\tau$-field. In the following, we consider $\sigma$ to 00PN order, and for simplicity, set $\sigma=0=\Sigma$ to $1$PN order. The meaning of $\sigma=0=\Sigma$ will be explained later.
@@
Later analysis shows that the nonvanishing ${\cal K}$ is 1PN order. To 0PN order the results are exactly the same as using the model-I in Eq. (9) with $n=1$. That is, only $n=1$ contributes to a mass-like term to 0PN order Blanchet-Skordis-2024 .
The observer’s four-vector $u_{a}$ to 1PN order is presented in Eq. (B7) of Hwang-Noh-axion-2023 ; to 0PN order we have $u_{i}=av_{i}/c$ where the index of $v_{i}$ is associated with $\delta_{ij}$ and its inverse. Using
@@
where indices of $Q_{i}$ and $\Pi_{ij}$ are associated with $\delta_{ij}$ and its inverse, the $\tau$-fluid quantities in Eq. (6) become
$\mu={1\over 8\pi G}\Big{\{}{\cal J}c^{4}-2c^{2}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}+\nu^{2}\Xi\Big{]}\Big{\}},$
$p=-{1\over 8\pi G}{\cal J}c^{4},$
$Q_{i}={2c^{2}\over 8\pi G}\Big{(}v_{i}+{1\over a}\sigma_{,i}\Big{)}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,j})_{,j}+\nu^{2}\Xi\Big{]},$
$\Pi_{ij}=0.$ (29)
To 0PN order, we have $\mu=\varrho c^{2}$. As we set ${\cal J}=0$ to the background order, the $\tau$-fluid quantities do not contribute to the background. As we have ${\cal J}=\delta{\cal J}={\cal J}_{,A}A$ and $A\sim{\cal O}(c^{-4})$, the first term in $\mu$ is higher PN order compared with the second term, and similarly the $p$ is higher PN order compared with $\mu$. Thus, we have
$\varrho=-{1\over 4\pi G}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}+\nu^{2}\Xi\Big{]}.$ (30)
@@
The $\tau$-fluid is coupled only to the longitudinal part of the velocity. The energy-momentum tensor of $\tau$-field becomes
$T^{0}_{0}=-\mu,\quad T^{0}_{i}=-c\varrho\sigma_{,i},\quad T^{i}_{0}={c\over a^{2}}\varrho\sigma^{,i},\quad T^{i}_{j}=p\delta^{i}_{j}.$ (32)
@@
${1\over a^{3}}(a^{3}S^{0})^{\displaystyle{\cdot}}-{1\over a^{2}}(S^{0}\sigma^{,i})_{,i}=0,$ (33)
$2\Phi_{,i}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,j})_{,j}+\nu^{2}\Xi\Big{]}=-{\cal J}_{,i}c^{4},$ (34)
where
$S^{0}=2\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}+\nu^{2}\Xi\Big{]}=-8\pi G\varrho,$ (35)
and $S^{i}=-S^{0}\sigma^{,i}/(ca^{2})$. Equation (33) is the same as equation of motion. In terms of $\tau$-fluid quantities, Eqs. (33) and (34) become
$\dot{\varrho}+3H\varrho-{1\over a^{2}}(\varrho\sigma^{,i})_{,i}=0,$ (36)
$\varrho\Phi_{,i}=-p_{,i}.$ (37)

### v2

@@ Abstract
A relativistic MOND theory, promising in reproducing cosmology as well as the MOND phenomenology in the low acceleration regime, was recently proposed. We present the post-Newtonian (PN) approximation and relativistic perturbation equations of this theory in cosmological context. The PN equations are presented to 1PN order and perturbation equations are presented in fully-nonlinear and exact forms. The gauge issues are clarified. The 1PN equations and linear perturbation equations are presented without imposing temporal gauge conditions. We show that to 0PN order baryon perturbation grows faster in the MOND regime. The MOND field can be interpreted as a fluid with specific equation of state without anisotropic stress, and the Jeans criterion is derived for the MOND field.
@@ I Introduction
Milgrom, in 1983, proposed MOND (Modified Newtonian Dynamics) to explain the gravity anomaly in galaxy rotations MOND-1983a . Instead of a popular way of saving conventional gravity (both Newton’s and Einstein’s) by introducing unseen dark matter (DM), he has modified in a non-relativistic context either gravity or inertia for the anomaly. Despite the lack of agreed relativistic theory, thus unable to address the cosmological demands other than the galactic rotation and related phenomena on similar scales, the non-relativistic proposal has survived till today as an alternative to the dark matter hypothesis. Recently, it gained independent observational support and challenge in stellar and Solar system scale phenomena; see Chae-2023 ; Hernandez-2023 ; Chae-2024a ; Chae-2024b ; Hernandez-2024 ; Kroupa-2024 for support from the long period binary stars observed by GAIA and from tidal tail studies of open clusters, and Vokrouhlicky-2024 ; Desmond-2024 for challenges to specific MOND models in explaining the gravitational field and the long period dynamics in the solar system.
@@
Observations have pin down the Milgrom’s universal constant of nature $a_{M}\simeq 1.2\times 10^{-8}cm/sec^{2}$ McGaugh-2004 . As originally noticed by Milgrom MOND-1983a the transition acceleration $a_{M}$ is intriguingly close to the acceleration scale of the expanding Universe $cH_{0}/(2\pi)\sim 1.1\times 10^{-8}cm/sec^{2}$ or $c^{2}\sqrt{\Lambda}=\sqrt{3\Omega_{\Lambda 0}}cH_{0}\sim 1.4cH_{0}$ using the current value of Lemaître-Hubble parameter $H_{0}\sim 70(km/sec)/Mpc$ and $\Omega_{\Lambda 0}\sim 0.69$; we have $cH_{0}=L_{H}/t_{H}^{2}$ with $L_{H}\equiv c/H_{0}$ and $t_{H}\equiv 1/H_{0}$. The interpolating function $f$ is not determined yet.
Bekenstein and Milgrom, in 1984, proposed a non-relativistic Lagrangian formulation of MOND, often called as AQUAL, leading to a modified Poisson’s equation $\nabla\cdot(f\nabla\Phi)=4\pi G\varrho$ with $\vec{a}=\nabla\Phi$ Bekenstein-Milgrom-1984 . At large distance from an isolated body with high symmetry, Milgrom’s original proposal is recovered Bekenstein-Milgrom-1984 . In the MOND based on the modified Poisson’s equation, the internal gravitational dynamics of a system is affected by an external gravitational field, often called the external-field effect, thus violating the strong equivalence principle Bekenstein-Milgrom-1984 . A variety of alternatives to AQUAL theory producing the MOND phenomenology is suggested Milgrom-2010-QUMOND ; Milgrom-2023-TriMOND ; Milgrom-2023-GQUMOND , and diverse relativistic extensions are also proposed Bekenstein-Milgrom-1984 ; Famaey-2012 .
Blanchet and Marsat Blanchet-Marsat-2011 introduced a vector field in Einstein’s gravity to recover the MOND phenomenology in weak gravity low acceleration regime. They used a function of the acceleration of a four-vector $U_{a}$ associated with a scalar function $\tau$ as $U_{a}\propto\nabla_{a}\tau$. The theory is motivated by a quantum gravity proposal with local Lorentz invariance violation Horava-2009 ; Blas-2010 and Einstein aether theories Jacobson-Mattingly-2001 ; Zlosnik-2007 ; Jacobson-2010 ; Sanders-2011 .
The MOND proposal is successful in explaining the phenomenology in the galactic scale without demanding DM. However, successful cosmology also demands the presence of a sort of DM (a fluid with dust-like equation of state) in explaining phenomena in the large-scale structure, the cosmic background radiation (CMB), and the background matter content. Blanchet and Skordis Blanchet-Skordis-2024 recently proposed one more additional function in the Blanchet and Marsat’s model, now using a function of the normalization amplitude of the same four-vector $U_{a}$ to accommodate such demands in cosmology.
@@
Here, we present equations for the post-Newtonian (PN) approximation and relativistic perturbation theory in cosmological context. The PN approximation is presented to 1PN order, and the perturbation theory is presented in fully-nonlinear and exact form. We address the gauge issues in both formulations and the linear perturbation equations are presented without imposing temporal gauge conditions. These subjects are already studied in Blanchet-Skordis-2024 and here we elaborate some aspects. For both formulations, we compare the MOND contribution with an ordinary fluid. To the linear order in perturbation, the MOND field does not contribute to the vector- and tensor-type perturbations, and it can be interpreted as a fluid with a specific equation of state (depending on the functional forms of ${\cal J}$ and ${\cal K}$) and is not associated with the anisotropic stress.
@@ II Relativistic MOND
$\hskip-22.76228ptU_{a}\equiv-{c\over Q}\nabla_{a}\tau,\quad Q\equiv c\sqrt{-(\nabla^{c}\tau)\nabla_{c}\tau},$
@@
and $L_{m}$ is the ordinary matter (including baryon and photon) part Lagrangian. Although we call the fundamental field simply a $\tau$-field, what is introduced is a four-vector field $U_{a}$ defined using the $\tau$-field; $U_{a}$ is a normalized timelike four-vector with $U^{c}U_{c}\equiv-1$; $q_{ab}\equiv g_{ab}+U_{a}U_{b}$ is a spatial projection tensor orthogonal to $U_{a}$; $A_{a}$ is an acceleration of the four-vector. For $\tau=\tau(t)$, we have $U_{a}=n_{a}$ where $n_{a}$ is the normal four-vector with $n_{i}\equiv 0$.
@@
The MOND phenomenology will constrain the form of ${\cal J}$. We will often consider two models for ${\cal K}$ Blanchet-Skordis-2024
@@ III Cosmological Background
where $x^{0}=\eta$ is the conformal time. Using $\tau=\tau(t)$, we have
@@
where an overdot indicates a derivative with respect to the cosmic time $t$. Equation (4), with $T_{00}=a^{2}\mu$ and $T_{ij}=a^{2}p\delta_{ij}$, gives $\tau$-fluid quantities
@@ III.1 Background constraint
$Q=1+\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n},\quad\mu_{\tau}={c^{4}I_{0}\over 8\pi Ga^{3}}{nQ+1\over n+1},$
$w_{\tau}={Q-1\over nQ+1},\quad c_{\tau}^{2}={1\over n}\Big{(}1-{1\over Q}\Big{)},$ (19)
where $a_{0}=a(t_{0})\equiv 1$ at present epoch. By a suitable parameter choice, with $[I_{0}/(2\nu^{2}{\cal K}_{n+1}a^{3})]^{1/n}\ll 1$, ${\cal K}$-term can simulate a dust fluid with $w_{\tau}\simeq 0\simeq c_{\tau}^{2}$ in recent era. As $Q$ approaches $1$ the $\tau$-field behaves like a zero-pressure fluid to the background order with $I_{0}=8\pi G\mu_{\tau 0}a_{0}^{3}/c^{4}$. For $Q\neq 1$, $Q$ increases as $a$ decreases. Far in the early era with $I_{0}/(2\nu^{2}{\cal K}_{n+1}a^{3})\gg 1$, we have $w_{\tau}\simeq c_{\tau}^{2}\simeq 1/n$ and $\mu_{\tau}\propto a^{-3(1+1/n)}$ Blanchet-Skordis-2024 .
For the dust-like behavior of $\tau$-field in the background, the authors of Blanchet-Skordis-2024 demand $w_{\tau*}\leq 0.0164$ at $a_{*}\sim 10^{-4.5}$. Using $I_{0}\simeq 8\pi G\mu_{\tau 0}/c^{4}=3\Omega_{\tau 0}H_{0}^{2}/c^{2}$, we have
@@
For $n=1$, with $\Omega_{\tau 0}=0.26$, the upper limit gives $L_{*}=220pc$. We have $L_{*}=62$, $22$, $9.4$, $4.6$, and $0.57pc$, respectively, for $n=2,3,4,5$, and $10$. This is a constraint on $\nu\sqrt{{\cal K}_{n+1}}$ parameter from the requirement of dust-like behavior of $\tau$-field in the background cosmology Blanchet-Skordis-2024 .
@@ IV PN approximation
where the index of $P_{i}$ is associated with $\delta_{ij}$ and its inverse. In this form we already imposed spatial gauge conditions Chandrasekhar-1965 ; Hwang-Noh-Puetzfeld-2008 . In the PN approximation, we use $x^{0}=ct$. To 0PN order, $\Phi$ is the Newtonian gravitational potential; $\Psi$ and $P_{i}$ parts are 1PN order. To 1PN order, $\Phi$ contains 1PN order contribution as well. In Chandrasekhar’s notation, we have Chandrasekhar-1965 ; Hwang-Noh-Puetzfeld-2008
@@
where $\sigma$ and $\Sigma$ are functions of spacetime corresponding to 0PN and 1PN order, respectively. As we take $\tau=t$ to the background order, we have $Q=1$ and $U_{a}=(-1,0,0,0)$, thus ${\cal K}=0$ and ${\cal J}=0$, and the $\tau$-field does not contribute to the background order equations. Thus, our PN formulation is valid in the Friedmann background without ${\cal K}$ contribution, see below. Eq. (48).
@@
Apparently, the expansion is lengthy if we keep perturbations of $\tau$-field. In the following, we consider $\sigma$ to 00PN order, and for simplicity, set $\sigma=0=\Sigma$ to $1$PN order. The meaning of $\sigma=0=\Sigma$ to 1PN order will be explained later.
@@
To 0PN order the results are exactly the same as using the model-I in Eq. (9) with $n=1$. That is, only $n=1$ contributes to a mass-like term in Eq. (38) to 0PN order Blanchet-Skordis-2024 .
The observer’s four-vector $u_{a}$ to 1PN order is presented in Eq. (B7) of Hwang-Noh-axion-2023 ; to 0PN order we have $u_{i}=av_{i}/c$ where the index of $v_{i}$ is associated with $\delta_{ij}$ and its inverse. We introduce
@@
where indices of $Q_{i}$ and $\Pi_{ij}$ are associated with $\delta_{ij}$ and its inverse. Using Eq. (4), Eq. (5) for the $\tau$-field gives
$T_{00}=-{2c^{2}\over 8\pi G}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}+\nu^{2}\Xi\Big{]},$
$T_{0i}=-{2c\over 8\pi G}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,j})_{,j}+\nu^{2}\Xi\Big{]}\sigma_{,i},$
$T_{ij}=-{2\over 8\pi G}\Big{\{}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,k})_{,k}+\nu^{2}\Xi\Big{]}\sigma_{,i}\sigma_{,j}$
$\qquad-{\cal J}_{,A}\Xi_{,i}\Xi_{,j}-{1\over 2}a^{2}(\nu^{2}\Xi^{2}-{\cal J}c^{4})\delta_{ij}\Big{\}},$ (29)
where ${\cal J}c^{4}$ is $c^{0}$-th order; as we set ${\cal J}=0$ to the background order, ${\cal J}=\delta{\cal J}={\cal J}_{,A}A$ and $A\sim{\cal O}(c^{-4})$. Using Eq. (6), we have fluid quantities associated with the fluid four-vector $u_{a}$ of the $\tau$-fluid
$\hskip-14.22636pt\mu=-{2c^{2}\over 8\pi G}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}+\nu^{2}\Xi\Big{]},$
$\hskip-14.22636ptp={1\over 8\pi G}{2\over 3}\Big{\{}{1\over a^{2}}{\cal J}_{,A}\Xi^{,i}\Xi_{,i}+{3\over 2}(\nu^{2}\Xi^{2}-{\cal J}c^{4})$
$\qquad\hskip-14.22636pt-\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,k})_{,k}+\nu^{2}\Xi\Big{]}\Big{(}v^{i}+{1\over a}\sigma^{,i}\Big{)}\Big{(}v_{i}+{1\over a}\sigma_{,i}\Big{)}\Big{\}},$
$\hskip-14.22636ptQ_{i}=-\mu\Big{(}v_{i}+{1\over a}\sigma_{,i}\Big{)},$
$\hskip-14.22636pt\Pi_{ij}={1\over 8\pi G}\Big{\{}{2\over a^{2}}{\cal J}_{,A}\Big{(}\Xi_{,i}\Xi_{,j}-{1\over 3}\Xi^{,k}\Xi_{,k}\delta_{ij}\Big{)}$
$\qquad\hskip-14.22636pt-2\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,k})_{,k}+\nu^{2}\Xi\Big{]}\Big{[}\Big{(}v_{i}+{1\over a}\sigma_{,i}\Big{)}\Big{(}v_{j}+{1\over a}\sigma_{,j}\Big{)}$
$\qquad\hskip-14.22636pt-{1\over 3}\Big{(}v^{k}+{1\over a}\sigma^{,k}\Big{)}\Big{(}v_{k}+{1\over a}\sigma_{,k}\Big{)}\delta_{ij}\Big{]}\Big{\}}.$ (30)
To 0PN order, we have $\mu=\varrho c^{2}$.
@@
The $\tau$-fluid is coupled only to the longitudinal part of the velocity. Fluid quantities become
$\varrho=-{2\over 8\pi G}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}+\nu^{2}\Xi\Big{]},$
$p={1\over 8\pi G}\Big{(}{2\over 3a^{2}}{\cal J}_{,A}\Xi^{,i}\Xi_{,i}+\nu^{2}\Xi^{2}-{\cal J}c^{4}\Big{)},$
$\Pi_{ij}={1\over 8\pi G}{2\over a^{2}}{\cal J}_{,A}\Big{(}\Xi_{,i}\Xi_{,j}-{1\over 3}\Xi^{,k}\Xi_{,k}\delta_{ij}\Big{)}.$ (32)
The energy-momentum tensor of $\tau$-field becomes
$T^{0}_{0}=-\varrho c^{2},\quad T^{0}_{i}=-\varrho c\sigma_{,i},\quad T^{i}_{0}={1\over a^{2}}\varrho c\sigma^{,i},$
$T^{i}_{j}=p\delta^{i}_{j}+\Pi^{i}_{j}+{1\over a^{2}}\varrho\sigma^{,i}\sigma_{,j}.$ (33)
@@
${1\over a^{3}}(a^{3}\varrho)^{\displaystyle{\cdot}}-{1\over a^{2}}(\varrho\sigma^{,i})_{,i}=0,$ (34)
$-\dot{\sigma}_{,i}+{1\over a^{2}}\sigma^{,j}\sigma_{,ij}+\Phi_{,i}+{1\over\varrho}(p_{,i}+\Pi^{j}_{i,j})=0.$ (35)
Equation (34) is the same as equation of motion with $S^{0}=-8\pi G\varrho$ and $S^{i}=-S^{0}\sigma^{,i}/(ca^{2})$. By using $v_{i}$, the fluid velocity of $\tau$-field, we have
${1\over a^{3}}(a^{3}\varrho)^{\displaystyle{\cdot}}+{1\over a}(\varrho v^{i})_{,i}=0,$ (36)
${1\over a}(av_{i})^{\displaystyle{\cdot}}+{1\over a}v^{j}v_{i,j}+{1\over a}\Phi_{,i}+{1\over\varrho a}(p_{,i}+\Pi^{j}_{i,j})=0.$ (37)
These are formally identical to conservation equations of an ordinary fluid with $p$ and $\Pi_{ij}$ given in Eq. (32); for the baryon, we have $p=0=\Pi_{ij}$. We can show $p_{,i}+\Pi^{j}_{i,j}=-\varrho\Xi_{,i}$, and using Eq. (25), Eq. (37) is identically valid for the $\tau$-field.